An object falls from rest, and s =-16t² where s feet is the distance of the object from the starting point at t seconds, and the positive direction is upward. If a stone is dropped from a building 256 ft high, find


A. the instantaneous velocity of the tone 1 sec after it is dropped;


B. the instantaneous velocity of the stone 2 sec after it is dropped;


C. how long it takes the stone to reach the ground;


D. the instantaneous velocity ot the stone when it reaches the ground.



1
Expert's answer
2021-03-07T17:26:06-0500

s=-16t2 g=32ft/sec2,

a) instantaneos velocit of th stone is v(t)=g*t=32*1=32ft/sec.

b) the same for this case also, v(2)=32*2=64ft/sec.

c)total distance travelled id S=1/2*g*t2=> from tis we can find t="\sqrt{2S/g}" ="\sqrt{2*256/32}" =4 sec.

d) v(4)=g*t=32*4=128ft/sec